Determine the liming requirement for a soil (acre furrow slice 6 inch depth) with the following properties:
Soil CEC: 28 cmolc/kg soil
Initial pH: 4.7
Initial %Base Saturation: 45%
Target pH: 6.5
Target % Base Saturation: 85%
Answer #1

Ans) Given,

CEC (Cation exchange capacity) = 28 cmol /kg or 28 meq /100 gm

Change in pH = 6.5 - 4.7 = 1.8

Change in base saturation = target saturation - initial saturation = 85 - 45 = 40 %

We know,1200 lb lime (CaCO3) per acre will require to change pH of about 1.8 per 100 gm

For change in base saturation by 40% , lime utilized = 0.4 x 28 meq/100 gm = 11.2 meq/100 gm

=> Lime required = 11.2 meq x 1200 lbs = 13440 lb  

1 ton = 2000 lb , so 13440 lb = 6.72 ton